OpenStudy (anonymous):

Enrique is 5 years younger than his sister, Josefina. Josefina is 15. How old is Enrique? Let E = Enrique's age.

OpenStudy (anonymous):

E = J - 5 with J for Josefina's age So E = 15-5 = 10